SSG Cycles manufactures and distributes motorcycle parts and supplies. Employees are offered a variety of share-based compensation plans. Under its nonqualified stock option plan, SSG granted options to key officers on January 1, 2021. The options permit holders to acquire 7 million of the company’s $1 par common shares for $27 within the next six years, but not before January 1, 2024 (the vesting date). The market price of the shares on the date of grant is $29 per share. The fair value of the 7 million options, estimated by an appropriate option pricing model, is $8.10 per option.

Required:

1. Determine the total compensation cost pertaining to the incentive stock option plan.
2. & 3. Prepare the appropriate journal entries to record compensation expense on December 31, 2021, 2022, and 2023. Record the exercise of the options if all of the options are exercised on May 11, 2025, when the market price is $30 per share.
Answer #1
Values Given in the Question:
Grant Date = 01/01/2021
Fair Value Per Option = $8.10 
Vesting Period = 6 years
Vesting date = 01/01/2024
Par Value of per Common share = $1
Market price of the shares on the date of grant= $29
Exercise price per option = $27
Exercise Date = 11/05/2025
Market price on Exercise date = $30 per share

Options expected to vest = 7000000
Calculation of total compensation cost:

Total Stock option compensation cost = Options x Fair value per option at grant date

= 7000000 x 8.10

= $56700000


Compensation Expense for 1st year [on 31 December 2021] :
Total stock option compensation = $56700000 
Vesting period = 6 years 
Service period completed = 1 year

Cumulative expense at end of year 1 = Total cost x Service period / Vesting period

Cumulative expense at end of year 1 = 56700000 x 1/6 = $9450000
Previously recognized expense = 0
Stock option compensation expense for year 1 = $9450000

Journal Entries – [on 31 December 2021] :

Stock option compensation expense Dr $9450000

To Stock Option Outstanding $9450000

Compensation Expense for 2nd year [on 31 December 2022] :
Total stock option compensation = $56700000 
Vesting period = 6 years 
Service period completed = 2 year

Cumulative expense at end of year 2 = Total cost x Service period / Vesting period

Cumulative expense at end of year 2 = 56700000 x 2/6 = $18900000
Previously recognized expense = $9450000

Hence Stock option compensation expense for year 2 = ($18900000-$9450000)

= $9450000

Journal Entries – [on 31 December 2022] :

Stock option compensation expense Dr $9450000

To Stock Option Outstanding $9450000

Compensation Expense for 3rd year [on 31 December 2023] :
Total stock option compensation = $56700000 
Vesting period = 6 years 
Service period completed = 3 year

Cumulative expense at end of year 3 = Total cost x Service period / Vesting period

Cumulative expense at end of year 3 = 56700000 x 3/6 = $28350000
Previously recognized expense = $18900000

Hence Stock option compensation expense for year 3 = ($28350000-$18900000)

= $9450000

Journal Entries – [on 31 December 2023] :

Stock option compensation expense Dr $9450000

To Stock Option Outstanding $9450000

Accounting Entry on 11 May 2025

We have to recognize the balance amount of compensation expense in this year

Stock option compensation expense Dr $9450000

To Stock Options Outstanding $9450000

Entry on Exercise;

Bank [7000000 x 27] Dr $189000000

Stock Options Outstanding Dr. $56700000

To Share Capital $7000000

To Securities Premium $238700000